Consider the following information: State of Economy Probability of State of Economy Rate of Rtn Stock A Rate of Rtn Stock B Rate of Rtn Stock C Boom .20 .24 .45 .33 Good .35 .09 .10 .15 Poor .30 .03 -.10 -.05 Bust .15 -.05 -.25 -.09 a. Your portfolio is invested 30 percent each in A and C, and 40 percent in B. What is the expected return of the portfolio? b. What is the variance of this portfolio? c. What is the standard deviation of this portfolio? d. What does the standard deviation reveal?
Answer #1

a) expected return of post folio: peso state of economy :- Boom stock proportion returen (a) (b) A - 0.3 0.24 B 0.4 0.45 c 0state of economy; Bust stork A B c propostion 6.3 0.4 0.3 expected return -0.05 -0.25 -Oog retuan expected return -0.015 - 0(6) Variance of post folio = 0.2 (0.351-0.0743)2 + 0.35 (0-112 - 0.0743) +0.30 (-0.046 - 0.0743)2 +0.15 (-0-142–0-074372 = 0.(d) standard deviation reveals the volatility of refugios of stock. It measures the amount that the return deviates from - me